OpenStudy (anonymous):

You have two events A and B such that P(A|B) = 0.4 and P(A) = 0.4. Based on this information, what can you conclude?

OpenStudy (anonymous):

since knowing B has occurred does not change the probability of A occurring, those events are "independent"

OpenStudy (anonymous):

in fact that is the definition of independence

OpenStudy (anonymous):

so Event A is not dependent on Event B. @satellite73

jimthompson5910 (jim_thompson5910):

event A alone P(A) = 0.4 the probability of event A occurring is 0.4 (ie 40%) now add in event B we are given event B happens because P(A|B) = 0.4, but this probability is identical to P(A) since P(A) = P(A|B), this means B does not change A so yes, A is not dependent on B. If it were, then B would change A.

OpenStudy (anonymous):

Thank you peeps.

OpenStudy (mathmate):

hmm... P(A|B) = conditional probability of A given B =P(A\(\cap\)B) / P(B)

OpenStudy (mathmate):

@aleale97 Can you confirm that the question says P(A|B) = 0.4 (conditional proability) and not \(P(A\cap B)=0.4\) (intersection)? Since we have \(P(A\cup B) = P(A)+P(B)-P(A\cap B)\) we get \(P(A\cap B) = P(A)+P(B)-P(A\cup B)\) Substituting values into definition of conditional probability, P(A)=0.4, P(A|B)=0.4 \(P(A|B) = \dfrac{P(A)+P(B)-P(A\cup B)}{P(B)}\) \( 0.4 = \dfrac{0.4+P(B)-P(A\cup B)}{P(B)}\) Solving for P(B) \(P(B)=\dfrac{P(A\cup B)-0.4}{0.6}\) I am short of conclusions though. All I can say is that B \(could\) be the sample space \(\Omega\), where P(B)=1, but this is one of the possible scenarios, and does not have to be the case.